OpenStudy (anonymous):

Use a calculator to find sin-1(-0.83). Express your answer in radians, rounding to the nearest hundredth.

OpenStudy (anonymous):

\[\sin^{-1} (-0.83)\]

OpenStudy (anonymous):

@abbychandler

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhhhhhh yea nvm i thought you meant something else. this is trig, right?

OpenStudy (anonymous):

yep @abbychandler

OpenStudy (anonymous):

you know anyone that does know trig? @abbychandler

OpenStudy (anonymous):

no sorry :(

OpenStudy (anonymous):

@robtobey can you help?

OpenStudy (anonymous):

\[\sin ^{-1}(-0.83)=-56.0987 \deg \]

OpenStudy (anonymous):

@robtobey how would you convert that to radians? can you teach me?

OpenStudy (anonymous):

-0.979108 radians using the Mathematica program.

OpenStudy (anonymous):

Don't see an obvious conversion to exact symbolic form.

OpenStudy (anonymous):

Lol thank you I just figured that out @ robtobey Can you help with other questions?

OpenStudy (anonymous):

@robtobey can you help with other questions?

OpenStudy (anonymous):

go ahead.

OpenStudy (anonymous):

Thank you! ok so Calculate \[\tan[\arcsin(\frac{ \sqrt{2} }{ 2 })]\]

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

yeah I got it, its 1